# Large Multimodal Models for Low-Resource Languages: A Survey

### v1

@@ 9 Conclusion and Future Work
Conclusion. Our survey has provided a comprehensive analysis of LMM-based approaches for LR languages, comprising 106 studies across 75 languages. We noted that vision-language combinations dominate the current research landscape (63% of surveyed works), but there is also an increasing trend toward incorporating video and speech in recent works. Additionally, we observed a particular concentration of research in South Asian languages (including Hindi, Bengali, Malayalam) and Southeast Asian languages (including Vietnamese, Javanese, Malay), followed by contributions in Middle Eastern languages (Persian, Arabic) and African languages (Hausa, Amharic), leaving 40 other languages each represented by a single study. However, the landscape of LMMs for LR languages has shown remarkable progress across multiple dimensions, from data creation and engineering to different fusion techniques and architectural innovations. The emergence of projects like HVG, SEACrowd, and BVG highlights the growing attention to creating high-quality multimodal resources for traditionally understudied languages, from simple parallel corpus creation to synthetic data generation. Both fusion techniques and visual enhancement approaches have matured significantly, with advances ranging from early fusion architectures to advanced visual disambiguation methods. Recent successes with models such as Qalam, LaVy, and Amharic LLaVA demonstrate that carefully designed multimodal strategies can effectively leverage limited resources, while adapting large-scale architectures for LR contexts.
Future work. We identify several promising directions for future research in the area of LMMs for LR languages. While current work has established a foundation in vision-language applications, there is a clear need to explore additional modality combinations, particularly incorporating audio and video modalities for LR languages, which could enable more robust and realistic language learning. Additionally, the development of comprehensive multimodal datasets remains a challenge, particularly for languages outside major language families. To address this limitation, research should focus on two areas: advancing synthetic data generation techniques, as demonstrated by recent works such as HVG, ELAICHI and Vintern-1B, and improving cross-lingual transfer methodologies, building upon frameworks such as XtremeCLIP and LowCLIP. These approaches show promise in overcoming the persistent challenge of limited training data availability for LR languages.
We also identify several advanced methods that have not been extensively explored in the context of multimodal learning for LR languages. These include: late fusion based on stacking, which learns to optimally combine outputs of individual models; tensor fusion, which captures complex cross-modal interactions through outer product and decomposition operations; kernel-based fusion, which maps data to high-dimensional spaces for non-linear integration of heterogeneous modalities; and graphical fusion, which leverages graph-based representations and neural architectures for multimodal data. Investigating the adaptation of these techniques to LR settings, while addressing challenges of data scarcity and computational efficiency, presents promising avenues for future research. We further recommend several strategies for improving fusion techniques in LR settings. First, researchers should consider adaptive fusion mechanisms that can dynamically adjust the contribution of each modality based on input quality and task requirements. Second, more efficient architectural designs are needed to make sophisticated fusion techniques accessible in resource-constrained environments. Third, the development of robust evaluation metrics specifically for multimodal fusion in LR contexts would help better understand the effectiveness of different approaches. Finally, taking inspiration from the success of MRF Chauhan et al. (2022), future work should explore more hybrid approaches that combine the strengths of different fusion strategies, while maintaining computational efficiency. These recommendations aim to address the unique challenges of multimodal fusion in LR languages, while maximizing the benefits of complementary information from different modalities.

### v2

@@ 10 Conclusion and Future Work
Conclusion. Our survey has provided a comprehensive analysis of LMM-based approaches for LR languages, comprising 117 studies across 96 languages. Vision-language combinations dominate the current research landscape (65% of surveyed works), with an increasing trend toward incorporating video and speech in recent works. We observed a concentration of research in South Asian languages (including Hindi, Bengali, Malayalam), Southeast Asian languages (Vietnamese, Javanese, Malay), Middle Eastern languages (Persian, Arabic) and African languages (Hausa, Amharic), while 42 other languages appear in only one study each.
The landscape of LMMs for LR languages has shown remarkable progress across multiple dimensions, from data creation to fusion techniques and architectural innovations. Projects like HVG, SEACrowd, and BVG highlight growing attention to creating high-quality multimodal resources for understudied languages. Recent successes with models such as Qalam, LaVy, and Amharic LLaVA [2024-andersland] demonstrate that carefully designed multimodal strategies can effectively leverage limited resources, while adapting large-scale architectures for low-resource contexts.
Challenges and gaps. Our analysis reveals several critical challenges in the current landscape of LMMs for LR languages. A significant modality imbalance exists, with text-image pairs dominating research (65% of studies), while audio and video modalities remain underexplored. This gap is particularly problematic for languages with strong oral traditions, where speech, tone and gesture carry essential linguistic information, with only 32% of studies incorporating audio, despite its crucial importance for predominantly oral languages, and only 8.5% of studies incorporating a video modality. We also identified persistent dataset scarcity and uneven language representation, with just three languages (Hindi, Arabic, Bengali) accounting for a disproportionate share of research attention. Technical limitations further constrain progress, as computational constraints limit the application of advanced fusion techniques in resource-constrained environments typical for LR contexts. Current cross-modal transfer methods struggle with catastrophic forgetting and inefficient knowledge transfer, particularly for languages that are structurally distant from high-resource counterparts. The field also lacks standardized evaluation frameworks for meaningful comparison across approaches, while recent work by Shen et al. [2024-shen] highlights significant safety challenges when deploying LLMs in multilingual contexts. Finally, sociolinguistic dimensions remain underexplored, including cultural representation, algorithmic bias, and potential impacts on language endangerment and revitalization efforts. These concerns are particularly acute given power imbalances between communities speaking low-resource languages and the primarily Western institutions developing these technologies.
Our study identifies three mechanisms through which LMMs may perpetuate digital inequalities. First, language model training inherently favors languages with larger training representation [navigli2023biases, wu2025inequalities], introducing a bias towards modeling HR languages. Second, benchmarks derived from English (e.g. translated MMLU) embed Western cultural assumptions that disadvantage LR language speakers even when linguistic accuracy is achieved [singh2024globalmmu], introducing cultural biases in the evaluation. Third, computational requirements exclude researchers in LR language regions from model development, creating dependency on external institutions and biasing resource access. Addressing these biases requires community-centered approaches that prioritize local capacity building, alongside technical performance metrics.
Future work. Based on the challenges identified above, we propose several key directions for future research.
For short-term development, we propose the following actionable research directions for benchmark and dataset creation: (1) extend Visual Genome-style multimodal datasets to at least 20 additional LR languages, prioritizing the 42 languages currently represented by only a single study; (2) develop speech-image paired corpora for tonal languages (e.g. Yoruba, Igbo, Fongbe), where audio modality carries critical semantic distinctions absent in text; and (3) establish a standardized “LR-MMBench” evaluation suite with culturally-adapted visual question answering tasks, following SEACrowd’s multilingual methodology, but incorporating non-Western visual contexts and evaluation protocols validated by native speakers.
To develop and improve LMMs for LR language, several concrete directions emerge from our analysis: (1) develop catastrophic forgetting mitigation strategies maintaining over 95% source-language performance, while achieving over 80% target-language performance for language pairs with fewer than 1,000 parallel sentences; (2) create language-agnostic visual encoders pre-trained on culturally-diverse image collections sourced from non-Western contexts, reducing the documented Western bias in current visual representations; and (3) establish explicit source-language selection guidelines based on typological similarity metrics (syntactic distance, shared writing systems, WALS features) to maximize positive transfer for specific target languages.
Several other research gaps require attention in future. Regarding the observed modality imbalance, researchers should prioritize incorporating audio and video for LR languages with limited writing traditions, enabling more robust applications that better reflect natural communication patterns, particularly for tonal languages and those where non-verbal communication is significant. For resource development, future work should advance synthetic data generation techniques (building on HVG, ELAICHI, Vintern-1B) and improve cross-lingual transfer methodologies (extending XtremeCLIP, LowCLIP) to accommodate greater linguistic diversity, while minimizing catastrophic forgetting. To overcome limitations in resource-constrained settings, researchers should investigate efficient fusion approaches including stacking-based late fusion, tensor fusion for complex interactions, and graphical fusion leveraging graph-based representations, all adapted for computational efficiency. Advancing adaptive integration through mechanisms that dynamically adjust the contribution of each modality based on input quality and task requirements will be crucial. Building on MRF [2022-chauhan], future work should explore hybrid approaches that combine strengths of different fusion strategies, while maintaining computational efficiency. Finally, adopting community-centered design approaches that address sociolinguistic dimensions alongside technical advances will ensure developments benefit the intended language communities themselves.
Finally, we advocate for mandatory community engagement through: (i) participatory design frameworks requiring documented language community involvement in dataset creation, with explicit data governance and benefit-sharing agreements; (ii) open-source, mobile-first data collection libraries suitable for field conditions, where many LR languages are spoken; and (3) standardized model cards for LR multimodal systems, documenting limitations, cultural biases, and appropriate use cases, ensuring transparent communication with end-user communities.
Acknowledgments
This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 351416. The authors thank reviewers for the constructive feedback.